OpenStudy (lgbasallote):

how to take the derivative of laplace transforms? like how to take the derivative of \[\mathcal L \{\sin kt\}\]

OpenStudy (anonymous):

L{sinkt} = k/s^2+k^2

OpenStudy (ash2326):

\[\frac d {ds} \int_0 ^{\infty} f(t) e^{-st}dt=\int_0 ^{\infty} -t\times f(t) e^{-st}dt=L[-t f(t)] \]

OpenStudy (anonymous):

u can go thru dat link

OpenStudy (lgbasallote):

yes i know what L {sin kt{ transforms to but i need to know how to derive it

OpenStudy (ash2326):

@lgbasallote you want to know how to derive L(sin kt)?

OpenStudy (anonymous):